# JEL classification: C11, C32, C55, E50

### v1

@@ 4 Conclusion
This paper presents a novel high-dimensional SVAR framework that allows for the imposition of a large number of linear inequality restrictions on both impact impulse responses and structural shocks, including their interactions. By leveraging a factor structure and introducing a computationally efficient sampling algorithm that avoids rejection-based procedures, our approach overcomes key limitations of existing methods, most notably their poor scalability in high dimensional settings. The flexibility to combine sign and shock restrictions enhances the precision of structural shock identification and allows for economically meaningful inference even in large systems.
Our empirical application, identifying five structural shocks in a 30 variable VAR, demonstrates the practical relevance of the methodology, particularly in uncovering the prominent role of financial shocks in driving business cycle fluctuations. As regards the narrative shock sign restrictions, we illustrate their capacity to sharpen inference by shrinking the identification set. More importantly, shock sign restrictions can facilitate the disentangling of structural shocks as shown for the financial risk and government risk shock. Overall, the proposed framework significantly broadens the scope for credible and computationally feasible structural analysis in large VAR environments.
@@
2001Q4: On November 14, 2001, amid a global economic slowdown intensified by the September 11 attacks, OPEC announced it would only cut oil production if non-OPEC producers did the same. This conditional stance was interpreted by markets as a potential signal of failed coordination and the risk of a price war, prompting a sharp downward revision in oil price expectations (Al-Naimi, (2016), of Oil Supply News: Evidence from OPEC Announcements, (2021)). We assume the shock to be negative.
2014Q4: A significant downward revision of oil price expectations occurred on November 27, 2014, when OPEC announced it would maintain existing production levels. Prior to the meeting, many market participants had anticipated that the cartel would implement production cuts to support falling oil prices. However, Saudi Arabia opposed calls from several economically weaker OPEC members to reduce output, resulting in a roughly 10 percent drop in price expectations (of Oil Supply News: Evidence from OPEC Announcements, , 2021). We assume the shock to be negative.

### v2

@@ 5 Conclusion
This paper presents a novel high-dimensional SVAR framework that accommodates a large number of linear inequality restrictions on both impact impulse responses and structural shocks, including their interactions. By incorporating a factor structure in the error terms and introducing a computationally efficient sampling algorithm that avoids rejection-based procedures, our approach addresses key limitations of existing methods - most notably, their limited scalability in high-dimensional settings. The flexibility to combine sign and shock restrictions enhances the precision of structural shock identification and enables economically meaningful inference even in large systems.
In a synthetic data exercise, we demonstrate the advantages of our proposed algorithm in large-scale VAR settings with multiple impact and shock sign restrictions, comparing our results to those of Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018). In our empirical application, we identify five structural shocks in a VAR with 30 variables, highlighting the practical utility of the method, particularly in revealing the prominent role of financial shocks in driving business cycle fluctuations. Regarding narrative shock sign restrictions, we show their effectiveness in sharpening inference by shrinking the identified set. More importantly, we demonstrate that shock sign restrictions can facilitate the disentanglement of structurally distinct shocks, as exemplified by the separation of financial risk and government risk shocks.
Overall, the proposed framework substantially broadens the scope for credible and computationally tractable structural analysis in high-dimensional VAR environments.
Disclosure statement
The authors have no conflicts of interest to declare. DeepL (version 1.47.0) was used to assist with grammar and language refinement in selected sections of the manuscript. All final edits and interpretations remain the sole responsibility of the authors. The data used in this study are available from the Federal Reserve Economic Data (FRED) database and from LSEG Data & Analytics.
@@
2001Q4: On November 14, 2001, amid a global economic slowdown intensified by the September 11 attacks, OPEC announced it would only cut oil production if non-OPEC producers did the same. This conditional stance was interpreted by markets as a potential signal of failed coordination and the risk of a price war, prompting a sharp downward revision in oil price expectations (Al-Naimi, (2016), Känzig, (2021)). We assume the shock to be negative.
2014Q4: A significant downward revision of oil price expectations occurred on November 27, 2014, when OPEC announced it would maintain existing production levels. Prior to the meeting, many market participants had anticipated that the cartel would implement production cuts to support falling oil prices. However, Saudi Arabia opposed calls from several economically weaker OPEC members to reduce output, resulting in a roughly 10 percent drop in price expectations (Känzig, , 2021). We assume the shock to be negative.
@@
Data Generation Process
In this Section, we describe the data generation procedure used to create synthetic datasets for our analysis in Section 3. We simulate a Vector Autoregressive (VAR) model with $n=10$ variables and $m=5$ structural shocks, using a sample size of $T=148$ observations and a lag length of $p=4$. The simulation includes a burn-in period of 100 observations ($T_{\text{full}}=T+\text{burnin}$) to ensure the model achieves stability before the actual data is then used for analysis.
We begin by generating random intercepts for each variable, drawn from a uniform distribution between 1-1- 1 and 111:
$\bm{b}_{0}\sim\mathcal{U}(-1,1),\$
with $\bm{b}_{0}$ being $n\times 1$.
The coefficients of the autoregressive matrix $B_{1}$ are drawn from uniform distributions:
$\bm{B}_{1}$$\sim$ $\mathcal{U}(0,0.3),\quad\text{for }i=j,$
$\bm{B}_{1}$$\sim$ $\mathcal{U}(-0.1,0.1),\quad\text{for }i\neq j.$
The coefficients of the autoregressive matrices for $p=2,4$ are drawn from normal distribution but are assumed to have smaller standard deviations which ensures that the coefficients decay over time:
$\bm{B}_{p,ij}\sim N(0,(0.05/\sqrt{p})^{2},\quad\text{for }p=2,\ldots,4\text{ and }i\text{ and }j.$
To ensure the VAR process is stable, we check the eigenvalues of the companion matrix $\bm{B}$. We assume that stability is achieved if the absolute values of all eigenvalues are less than 0.95:
$\bm{B}=\left[\begin{array}[]{cccc}B_{1}&B_{2}&B_{3}&B_{4}\\ I_{k}&0&0&0\\ 0&I_{k}&0&0\\ 0&0&I_{k}&0\end{array}\right]$ (29)
The stricter threshold condition of <0.95<0.95< 0.95 guarantees the stability of the VAR model.
The impact matrix $\bm{L}$ governs the relationship between the structural shocks and the variables in the system. We generate $\bm{L}$ from a normal distribution:
$\bm{L}\sim N(0,1).$
We ensure that each row is normalized to control the overall magnitude of the shocks:
$\bm{L}(i,:)=\frac{\bm{L}(i,:)}{\|\bm{L}(i,:)\|},\quad\text{for }i=1,\ldots,n.$
The structural shocks are drawn from a normal distribution with a smaller variance to ensure that the shocks have a moderate impact on the system:
$\bm{F}\sim\mathcal{N}(0,0.02),$
with $\bm{U}=\bm{F}\times\bm{L}^{\prime}$. This results in a $T_{\text{full}}\times m$ matrix of structural shocks applied to the system.
The simulated data is then generated recursively, with the following equation for each time step $t$:
$Y_{t}=\bm{b}_{0}+\bm{B}_{1}Y_{vect,t}+U+\mu+\sigma\cdot\epsilon_{t},$ (30)
where $Y_{vect,t}=vec([\bm{y}_{t-1},\bm{y}_{t-2},\ldots,\bm{y}_{t-p}])$, $\mu=1$ and $\sigma=0.5$.
After generating the full sample, we discard the burn-in period of 100 observations, retaining only the final 148 observations for analysis:
$Y=Y_{\text{full}}[\text{burnin}+1:T_{\text{full}}].$ (31)
This procedure results in a synthetic datasets with a controlled structure, which we use to evaluate the performance of our proposed algorithm and compare it with the approach by Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018).
Impact and Shock Sign Restrictions
As discussed in Section 3, we impose impact sign restrictions on $\bm{B}$ and shock sign restrictions on $\bm{F}$. For the impact sign restrictions, we randomly assign a 111 or 1-1- 1 to distinct positions in the $\bm{B}$, subject to two conditions: (i) each shock column must feature at least two sign restrictions, and (ii) the sign pattern of each shock must be unique, up to sign reversal (i.e., no two columns are identical or exact negatives). The procedure ensures that the resulting shock identification is both informative and non-redundant. The shock sign restrictions for $\bm{F}$ are directly derived from the signs of the sampled structural shocks $\bm{F}$. To establish the sign restrictions, time-shock pairs are randomly selected, and the corresponding signs of $\bm{F}$ are assigned to a shock sign restrictions matrix. In a robustness exercise that relaxed this condition, the algorithm proposed by Antol´ın-D´ıaz and Rubio-Ram´ırez, (2018) required substantially more time to generate an admissible draw that satisfied the shock sign restrictions.